Find all the solutions of the equation Sin x/2=1/2
Answer by lwsshak3(11628)

Find all the solutions of the equation Sin x/2=1/2
**
Sin x/2=1/2
Angles: π/6 and 5π/6 in quadrants I and II where sin>0
x/2=π/6+2πn+5π/6+2πn, n=integer
x=π/3+2πn+5π/3+2πn, n=integer